Speakers:
Claudio Luci
(Sapienza Universita e INFN, Roma I (IT)),
Daniela Katherinne Paredes Hernandez
(University of Hong Kong (HK)),
Enrico Pastori
(INFN e Universita Roma Tor Vergata (IT)),
Enrico Tusi
(INFN e Universita Roma Tor Vergata (IT)),
Eric Ballabene
(University and INFN, Bologna (IT)), Dr
Federico Lasagni Manghi
(Universita e INFN, Bologna (IT)),
Giulio Aielli
(INFN e Universita Roma Tor Vergata (IT)), Mr
Mohamed Zaazoua
(University of Science and Technology of China (CN)),
Sinem Simsek
(Istinye University (TR))
- Vth tuning in noisy trigger tower s55_t4 on 10/09 [e-log]
BML5A13.CO.Ly0.ETA 1.0 → 0.8 V
BML5A13.CO.Ly1.ETA 1.0 → 0.0 V
The initial noise rate of 10.3 kHz has been lowered to 6.4 kHz.
- One CAEN Board went into unplugged status, affecting 4 Vpad channels [e-log]
- However, they are working from the trigger side.
- Cavern intervention during access on 12/09 [e-log]
- The ADC board Y_2904_X7_C2/H1/br02/crate0/slot08 that stopped working → power cycled.
- Problematic Vee channel for chamber BMS4C16.CO that dropped from 3.1 to 1.7 A [e-log] → fixed.
- One ADC board stopped working on 14/09 [e-log]
- Same ADC board that was recovered in the cavern access.